### v4

# Smoothing Variances Across Time: Adaptive Stochastic Volatility

Abstract
We introduce a novel Bayesian framework for estimating time-varying volatility by extending the Random Walk Stochastic Volatility (RWSV) model with Dynamic Shrinkage Processes (DSP) in log-variances. Unlike the classical Stochastic Volatility (SV) or GARCH-type models with restrictive parametric stationarity assumptions, our proposed Adaptive Stochastic Volatility (ASV) model provides smooth yet dynamically adaptive estimates of evolving volatility and its uncertainty. We further enhance the model by incorporating a nugget effect, allowing it to flexibly capture small-scale variability while preserving smoothness elsewhere. We derive the theoretical properties of the global-local shrinkage prior DSP. Through simulation studies, we show that ASV exhibits remarkable misspecification resilience and low prediction error across various data-generating processes. Furthermore, ASV’s capacity to yield locally smooth and interpretable estimates facilitates a clearer understanding of the underlying patterns and trends in volatility. As an extension, we develop the Bayesian Trend Filter with ASV (BTF-ASV) which allows joint modeling of the mean and volatility with abrupt changes. Finally, our proposed models are applied to time series data from finance, econometrics, and environmental science, highlighting their flexibility and broad applicability.
Keywords: Stochastic Volatility; Bayesian Smoothing; Bayesian State-Space Model; Global-Local Shrinkage Prior
1 Introduction
Volatility of a time series quantifies deviations from the mean and is fundamental to understanding the underlying data-generating process. In finance, volatility estimation is essential for pricing, risk assessment, and asset management (Bollerslev, 1986; Hull and White, 1987). In epidemiology, it aids in early outbreak detection and forecasting new cases of diseases (Sarkar and Chatterjee, 2017; Achcar et al., 2020; Kostoulas et al., 2021). Climate science measures volatility to study phenomena like tornadoes, droughts, and rainfall (Tippett, 2014; Modarres and Ouarda, 2014; Mehdizadeh et al., 2017). In engineering, volatility helps predict mechanical failures (Pham and Yang, 2010; Ma et al., 2017), while in hydrology, it informs understanding of streamflow variations (Otache, 2012; Wang et al., 2005, 2023a, 2023b).
(a) S&P 500 Index
(b) Electricity
(c) Bike Rentals
(d) Weekly Log Returns
(e) Electricity (mean-centered)
(f) Bike Rentals (mean-centered)
Figure 1: Figures 1(a), 1(b) and 1(c) display the S&P 500 index from 2012-01-01 to 2021-12-31, daily total electricity demand (in gigawatt hours) in New York State from 2015-07-01 to 2024-06-30, and daily bike rental counts via Capital Bikeshare by non-members in Washington metropolitan area from 2012-01-01 to 2017-12-31, respectively. Figures 1(d), 1(e) and 1(f) show the corresponding mean-centered series: weekly log returns for the S&P 500 index with the average log return subtracted, and series for electricity demand and bike rentals mean-centered using smoothing spline, respectively. The S&P 500 index data are obtained using the quantmod package in R (Ryan and Ulrich, 2024) from Yahoo Finance (Finance, 2025), electricity demand data from the New York Independent System Operator (NYISO) (EIA, 2024), and bike rental data from Capital Bikeshare (Bikeshare, 2024).
Classical volatility models include Autoregressive Conditional Heteroskedasticity (ARCH) (Engle, 1982), Generalized ARCH (GARCH) (Bollerslev, 1986), and Stochastic Volatility (SV) (Hull and White, 1987; Taylor, 2008; Melino and Turnbull, 1990). They all make the stationarity assumption that the unconditional variance is constant. However, a growing body of econometric literature points to evolving volatility patterns in stock markets around the world (Chou, 1988; French et al., 1987; Poon and Taylor, 1992; So et al., 1997; Su and Wang, 2020; Nishino and Kakamu, 2015). Moreover, the stationarity assumption is often unrealistic in fields like climate science and epidemiology, due to climate change and seasonal or outbreak-driven dynamics (Tippett, 2014; Kostoulas et al., 2021).
Figure 1 illustrates evolving volatility patterns in three real-world datasets: weekly log-returns of the S&P 500 index, daily electricity demand in New York State, and daily bike rental counts via Capital Bikeshare. All three exhibit nonstationary volatility patterns. In Figures 1(a) and 1(d), persistent periods of high and low volatility reflect changing market conditions. In Figures 1(b), 1(e), 1(c) and 1(f), both electricity and bike data show seasonal cycles in both mean and volatility, with elevated volatility during high-demand periods.
Existing approaches address such patterns by incorporating hidden Markov switching (HMS) structure into existing SV and GARCH models (Brunetti et al., 2008; Ardia et al., 2018; Hamilton and Susmel, 1994; Cai, 1994; Bauwens et al., 2010; Gray, 1996a; So et al., 1998; Hwang et al., 2004). However, HMS models suffer from three major limitations: they (1) require specification of the number of regimes, which is typically unknown; (2) are computationally intensive as pointed out in Gray (1996b) and Billio et al. (2016); and (3) impose a discrete regime-switching structure, which may not capture ever evolving volatility dynamics. For example, Figure 1(f) reveals a seasonal increase in volatility for bike rentals, with the magnitude of these fluctuations growing over time. For example, volatility in 2012 is substantially lower than in 2017. Such a pattern is difficult to reconcile with a model based on a fixed number of regimes, as it reflects a continuous evolution of volatility rather than discrete shifts.
We introduce the Adaptive Stochastic Volatility (ASV) model as an extension of the Random Walk Stochastic Volatility (RWSV) model (Ruiz, 1994; Harvey et al., 1994). RWSV assumes the log-variance of the observed process follows a Gaussian random walk with a constant variance. In contrast, ASV introduces two significant modifications to RWSV: it incorporates a time-varying variance for the log-variance increments and adopts a global-local shrinkage prior for adaptability. For the shrinkage prior, we specifically highlight a class of prior distributions called Dynamic Shrinkage Processes (DSP) by Kowal et al. (2019), as it provides a flexible framework for adaptive shrinkage for time series. To further improve robustness, we also consider a nugget variant of ASV, which adds an additional Gaussian noise to capture local fluctuations not explained by the smooth latent process.
The standout feature of our proposed model is its local adaptability, effectively estimating volatility in the presence of both gradual and abrupt changes in the volatility process with notable smoothness in between. It allows robustness against model misspecification, resulting in low prediction error across diverse data-generating processes (DGPs). In addition, ASV’s smooth volatility estimates provide clearer insights into the underlying patterns and trends. The model also naturally highlights periods of heightened volatility variation, offering additional interpretability in identifying structural shifts.
The remainder of this paper is organized as follows. Section 2 first introduces DSP and its theoretical properties, followed by a description of our proposed model (ASV) and its associated parameters. Section 3 provides an overview of the data augmentation techniques for efficient Gibbs sampling scheme; full conditional distributions are detailed in the supplementary material. Simulation study results are summarized in Section 4. Section 5 presents empirical analyses on three real-world datasets: the S&P 500 index, daily electricity usage in New York, and bike rental counts. In Section 6, the Bayesian Trend Filter with ASV (BTF-ASV), which jointly estimates the time-varying means and the variances, is proposed as an extension and applied to the sunspot number data.
2 Methodology
2.1 The Random Walk Stochastic Volatility Model
ASV may be viewed as an extension of RWSV (Ruiz, 1994; Harvey et al., 1994). Consider a zero mean process with $T$ observations, $\{y_{t}\}_{t=1}^{T}$, and its log-variance term $\{h_{t}\}_{t=1}^{T}$. SV assumes the observed process $y_{t}$ to follow a normal distribution with its log-variance term, $h_{t}$, following the lag order 1 autoregression. RWSV is a special case of SV in which the autoregressive coefficient is fixed at 1. Thus, the model is defined as:
$y_{t}=\exp\{h_{t}/2\}\epsilon_{t}$ $[\epsilon_{t}]\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1),$ (1)
$[\Delta h_{t}|\sigma^{2}_{h}]\sim N(0,\sigma^{2}_{h})$ $[\sigma^{2}_{h}]\sim\pi(\sigma^{2}_{h}).$
In RWSV, $h_{t}$ is governed by a time-invariant variance term $\sigma_{h}^{2}$, which determines the degree of variation between successive log-variances. A large $\sigma_{h}^{2}$ implies high probability of significant changes in the $h_{t}$ process, whereas a small $\sigma_{h}^{2}$ indicates high probability of minimal to no changes in $h_{t}$. In a frequentist framework, $\sigma_{h}$ is treated as a fixed parameter, often estimated via quasi-maximum likelihood (Ruiz, 1994; Harvey et al., 1994). A more recent study by Nishino and Kakamu (2015) employs a Bayesian framework for estimation with a non-informative inverse-gamma priors for $\pi(\sigma^{2}_{h})$.
2.2 Stochastic Volatility with Local Only Shrinkage
RWSV may fit poorly when the time series exhibits both significant and minimal changes, as it assumes a time-invariant innovation variance $\sigma^{2}_{h}$. Ideally, we would prefer $\sigma^{2}_{h}$ to adapt over time, remaining small during stable periods and increasing during periods of abrupt changes in $h_{t}$. A natural approach to address this limitation is to introduce sparsity in the variance process. Under the frequentist framework, sparsity in $\Delta h_{t}$ may be attained by imposing a penalty, such as the $\ell_{1}$ penalty used in LASSO (Tibshirani, 1996). In a Bayesian setting, near sparsity can be imposed via heavy-tailed priors on the variance parameter. Motivated by the Bayesian LASSO (Park and Casella, 2008) and the scale-mixture Gaussian representation of the Laplace distribution (West, 1987), we propose an intermediate model, RWSV with Bayesian LASSO (RWSV-BL):
$y_{t}=\exp\{h_{t}\}\epsilon_{t}$ $[\epsilon_{t}]\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1),$ (2)
$[\Delta h_{t}|\sigma^{2}_{h,t}]\sim N(0,\sigma^{2}_{h,t}|\Lambda)$ $[\sigma^{2}_{h,t}]\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Exp}\{(2\Lambda^{2})^{-1}\},$
where $\Lambda>0$ is a hyperparameter controlling the prior concentration. Following Park and Casella (2008), we place a Gamma prior on $\Lambda^{2}$ for full Bayesian inference. Under RWSV, $\sigma^{2}_{h,t}$ induces shrinkage by encouraging most increments to be small, while still allowing occasional large deviations.
2.3 Dynamic Shrinkage Processes and their Properties
Before introducing the proposed model, we first describe a class of global-local shrinkage priors known as DSP, a central component in our formulation. Global-local shrinkage priors have been widely studied for Gaussian observations, particularly in high-dimensional regression. Notable examples include the horseshoe prior (Carvalho et al., 2010), the horseshoe+ prior (Bhadra et al., 2015), and the triple gamma prior (Cadonna et al., 2019). Among these, DSP offers a particularly flexible formulation that generalizes several existing priors and naturally incorporates temporal dependence.
DSP has shown versatility across various applications. For example, Wu et al. (2024a) integrate DSP into a Bayesian dynamic linear model to estimate change points and score outliers. Schafer and Matteson (2024) apply it to construct the negative binomial Bayesian trend filter (NB-BTF) for adaptively smoothing integer-valued time series. Additionally, Wu et al. (2024b) combine the Bayesian trend filter with DSP and a machine learning-based regularization method to effectively distinguish micro-level drifts from macro-level shifts.
DSP serves as a prior for time-varying log-variance parameters. Define $v_{t}:=\log(\sigma^{2}_{h,t})$, where $v_{t}|\mu,\phi\sim DSP(a,b,\mu,\phi)$. DSP imposes an autoregressive structure on $\{v_{t}\}_{t=1}^{T}$, allowing for locally adaptive shrinkage through heavy-tailed innovations around a persistent mean process:
$v_{t}:=\log(\sigma^{2}_{h,t})=\mu+\phi(v_{t-1}-\mu)+\eta_{t}$ $[\eta_{t}]\stackrel{{\scriptstyle iid}}{{\sim}}Z(a,b,0,1),$ (3)
where $Z$-distribution has the following density function (Barndorff-Nielsen et al., 1982):
$f(z|a,b,\mu_{z},\sigma_{z})=(\sigma_{z}\mathrm{B}(a,b))^{-1}\bigg{(}\exp\bigg{\{}\frac{z-\mu_{z}}{\sigma_{z}}\bigg{\}}\bigg{)}^{a}\bigg{(}1+\exp\bigg{\{}\frac{z-\mu_{z}}{\sigma_{z}}\bigg{)}\bigg{\}}^{-(a+b)}.$
For fully Bayesian inference, we place weakly informative priors on $\mu$ and $\phi$, and describe them in detail in Section 3. Under this framework, the global parameter is defined as $\tau=\exp\{\mu(1-\phi)/2\}$ and the local shrinkage parameter is defined as $\lambda_{t}=\exp\{(\phi v_{t-1}+\eta_{t})/2\}$. The autoregressive coefficient $|\phi|<1$ controls temporal dependence in local shrinkage. When $\phi$ is close to 1, the process exhibits strong positive dependence: strong shrinkage tends to be followed by continued strong shrinkage, and weak shrinkage by weak shrinkage. When $\phi$ is close to -1, the process favors alternation between strong and weak shrinkage across time.
The $Z$-distribution on the local scale parameter $\eta_{t}$ governs the degree of local shrinkage at each time $t$. In our simulation study (Section 4) and empirical study (Section 5), we consider two specific instances of DSP. When $a=b=1/2$ and the autoregressive coefficient $\phi=0$, the prior reduces to the well-known horseshoe prior (HS) (Carvalho et al., 2010), since $\lambda_{t}=\exp\{\eta_{t}/2\}\sim C^{+}(0,1)$. When $\phi$ is estimated from the data with $a=b=1/2$, the prior induces temporal dependence in shrinkage and is referred to as the dynamic horseshoe (DHS) prior (Kowal et al., 2019). Other special cases of DSP include the Strawderman–Berger prior when $a=1/2$ and $b=1$ (Strawderman, 1971), the normal–exponential–gamma prior when $a=1$ and $b>2$ (Griffin and Brown, 2005), and the Jeffreys prior with $a=b\rightarrow 0$ (Figueiredo, 2003).
While DSP has an appealing formulation and a clear connection to well-known priors, its theoretical properties remain incomplete. In particular, DSP distinguishes itself from other global-local shrinkage priors by introducing temporal dependence in the local shrinkage parameter $\lambda_{t}$. Due to the time dependence, the original work by Kowal et al. (2019) focuses on the conditional behavior while the properties of the marginal distribution remain unexamined. We derive new results on the stationary distribution of DSP.
Define $z_{t,h}:=\phi^{h}\eta_{t-h}$. Infinite-order moving average representation of $v_{t}$ results in $v_{t}=\mu+\phi(v_{t-1}-\mu)+\eta_{t}=\mu+\sum_{h=0}^{\infty}z_{t,h}$, with $\eta_{t}\stackrel{{\scriptstyle iid}}{{\sim}}Z(1/2,1/2,0,1)$. As shown in Barndorff-Nielsen et al. (1982), $z_{t,h}$ is a scaled hyperbolic secant distribution, with $\mathbb{E}(z_{t,h}|\phi)=0$ and $\mathrm{Var}(z_{t,h}|\phi)=(|\phi|^{h}\pi)^{2}$. Consider $z_{t}:=\sum_{h=0}^{\infty}z_{t,h}.$
Theorem 1.
$z_{t}$ converges almost surely if and only if $|\phi|<1.$
Proof.
This directly follows from Kolmogorov’s three series theorem. Note that the variance $\sum_{h=0}^{\infty}\mathrm{Var}(z_{t,h}|\phi)$ converges if and only if $|\phi|<1$. Exact derivation is explored in Appendix A.1. ∎
Corollary 1.1.
When $|\phi|<1$,
$\mathbb{E}(z_{t}|\phi)=\mathbb{E}(\sum_{h=0}^{\infty}z_{t,h}|\phi)=\sum_{h=0}^{\infty}\mathbb{E}(z_{t,h}|\phi)=0.$
$\mathrm{Var}(z_{t}|\phi)=\mathrm{Var}(\sum_{h=0}^{\infty}z_{h}|\phi)=\sum_{h=0}^{\infty}\mathrm{Var}(z_{h}|\phi)=\sum_{h=0}^{\infty}\pi^{2}\phi^{2h}=\frac{\pi^{2}}{1-\phi^{2}}.$
Characteristic function of $z_{h,t}$ is $\mathrm{sech}(\pi\phi^{h}t)$. Thus, the characteristic function, $g(t)$, for $\sum_{h=0}^{\infty}z_{h,t},$ would be an infinite product of the characteristic function of $z_{h,t}$.
$g(t)=\prod_{h=0}^{\infty}sech(\pi\phi^{h}t),$ $-\frac{1}{2}<t<\frac{1}{2}.$
Similarly, for the moment generating function $M(t)$
$M(t)=\prod_{h=0}^{\infty}\mathrm{sec}(\pi\phi^{h}t),$ $-\frac{1}{2}<t<\frac{1}{2}.$
Theorem 2.
If $|\phi|=0.5$, $\eta_{t-h}\stackrel{{\scriptstyle iid}}{{\sim}}Z(1/2,1/2,0,1)$, and $z_{t,h}:=\phi^{h}\eta_{t-h}$, then $z_{t}:=\sum_{h=0}^{\infty}z_{t,h}\stackrel{{\scriptstyle a.s}}{{\to}}Logistic(0,2)$.
Proof.
Derivations are explored in Appendix A.2. ∎
Thus, assuming $\mu=0$ and $\phi=0.5$ for simplicity, we have the following density function for $v_{t}$, $\lambda_{t}:=exp\{v_{t}/2\}$ and $\kappa_{t}:=\frac{1}{1+exp\{v_{t}\}}$ for the stationary distribution of DSP:
$f(v_{t})=\frac{1}{8}\mathrm{sech}^{2}\bigg{(}\frac{v_{t}}{4}\bigg{)},$
$f(\lambda_{t})=\frac{1}{4\lambda_{t}}\mathrm{sech}^{2}\bigg{(}\frac{\log(\lambda_{t})}{2}\bigg{)}=\frac{1}{(1+\lambda_{t})^{2}},$
$f(\kappa_{t})=\frac{1}{8(\kappa_{t})(1-\kappa_{t})}\mathrm{sech}^{2}\bigg{(}\frac{1}{4}\bigg{(}\log(\frac{1-\kappa_{t}}{\kappa_{t}})\bigg{)}\bigg{)}=\frac{1}{\sqrt{\kappa_{t}(1-\kappa_{t})}}\frac{1}{2\kappa_{t}(1+\sqrt{\frac{1-\kappa_{t}}{\kappa_{t}}})^{2}}.$
(a) Prior densities on $\lambda_{t}$
(b) Prior densities on $\kappa_{t}$
Figure 2: Prior densities of $\lambda_{t}$ and $\kappa_{t}$ under the horseshoe prior and the stationary distribution of Dynamic Shrinkage Process (DSP) with $\eta_{t}\stackrel{{\scriptstyle iid}}{{\sim}}Z(1/2,1/2,0,1)$, $\phi=1/2$ and $\mu=0$.
Figure 2 compares the prior distribution of $\lambda_{t}$ and $\kappa_{t}$ for the horseshoe and DSP. The horseshoe prior in Carvalho et al. (2010) assumes $\lambda_{t}\sim C^{+}(0,1)$ and $\kappa_{t}\sim Beta(1/2,1/2)$. The density function of $g(\lambda_{t})$ and $g(\kappa_{t})$ under the horseshoe are:
$g(\lambda_{t})=\frac{2}{\pi(1+\lambda_{t}^{2})},$ $g(\kappa_{t})=\frac{1}{\pi\sqrt{\kappa_{t}(1-\kappa_{t})}}.$
For $\lambda_{t}$, the density function for DSP has heavier tails and a sharper spike at zero. Specifically, $f(\lambda_{t})>g(\lambda_{t})$ when $\lambda_{t}\in(0,\frac{2-\sqrt{(4-\pi)\pi}}{(\pi-2)})\cup(\frac{2+\sqrt{(4-\pi)\pi}}{(\pi-2)},\infty)$. In terms of $\kappa_{t}$, the DSP prior puts more weight around 0 and 1 than the horseshoe. This is due to the extra U-shaped term in $\kappa_{t}$, $1/(2\kappa_{t}(1+\sqrt{\frac{1-\kappa_{t}}{\kappa_{t}}})^{2})$, pushing the mass of the distribution to near 0 and 1.
Finally, the marginal prior on $\Delta h$ is considered:
Theorem 3.
Let $|\phi|=0.5$, $\eta_{t-h}\stackrel{{\scriptstyle iid}}{{\sim}}Z(1/2,1/2,0,1)$,$z_{t,h}:=\phi^{h}\eta_{t-h}$, $v_{t}:=\sum_{h=0}^{\infty}z_{t,h}$, $\lambda_{t}:=exp\{v_{t}/2\}$ and $\Delta h_{t}\sim N(0,\lambda_{t})$,
$\lim_{\Delta h_{t}\to 0}f(\Delta h_{t})=\infty,$
$K_{L}\log(1+\frac{4}{\Delta^{2}h_{t}})<f(\Delta h_{t})<K_{U}\log(1+\frac{2}{\Delta^{2}h_{t}}),$ $|\Delta h_{t}|>0$
where $K_{U}=\frac{1}{2\sqrt{2\pi}}$ and $K_{L}=\frac{1}{8\sqrt{2\pi}}$
Proof.
The proof is similar to Theorem 1 in Carvalho et al. (2010). We use the fact that $\forall x>0$,
$\frac{1}{2(1+x^{2})}\leq\frac{1}{(1+x)^{2}}\leq\frac{1}{(1+x^{2})}.$
Derivation is detailed in Appendix A.3. ∎
The bounds for the marginal distribution of DSP are similar to the ones from the horseshoe with the only difference being the constant factor $K_{L}$ and $K_{U}$. Under the horseshoe prior, $K_{U}=\frac{1}{\sqrt{2\pi^{3}}}$ and $K_{L}=\frac{1}{2\sqrt{2\pi^{3}}}$. Like horseshoe prior, DSP is also unbounded near the origin, which leads to super-efficiency in a sparse setting as shown in Carvalho et al. (2010).
2.4 Adaptive Stochastic Volatility via Global-Local Shrinkage
We propose ASV, which uses DSP explored in Section 2.3, to adaptively estimate time-varying volatility in a smooth yet locally adaptive manner. In ASV, the prior models the log-variance of the $k$th-order differences of $h_{t}$: $\Delta^{k}h_{t}|\sigma^{2}_{h,t}\sim N(0,\sigma^{2}_{h,t})$. As previously defined, $v_{t}:=\log(\sigma^{2}_{h,t})$ follows an autoregressive structure described in Equation 3. Unlike RWSV-BL, ASV incorporates both the global parameter $\mu(1-\phi)$ and the local parameter $\phi v_{t-1}+\eta_{t}$. The global parameter promotes overall smoothness, while the local parameter preserves local adaptivity. The resulting model, ASV, has the following hierarchical representation:
$y_{t}=\exp\{h_{t}\}\epsilon_{t}$ $[\epsilon_{t}]\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1),$ (4)
$[\Delta^{k}h_{t}|\sigma^{2}_{h,t}]\sim N(0,\sigma_{h,t}^{2})$ $[\log(\sigma^{2}_{h,t})|\mu,\phi]\sim DSP(a,b,\mu,\phi).$
The differencing order $k$ plays a central role in controlling the model’s smoothness. ASV can be viewed as a variance-process analogue of smoothing methods for the mean function, such as the Bayesian Trend Filter by Roualdes (2015) and Kowal et al. (2019), Hodrick-Prescott filter by Hodrick and Prescott (1997), $\ell_{1}$-trend filter by Kim et al. (2009) and Tibshirani (2014), and a smoothing spline. Penalizing the $k$th-order difference encourages the underlying signal to be locally well-approximated by a polynomial of degree $k-1$: $k=1$ promotes piecewise constant behavior, $k=2$ encourages piecewise linearity, and higher-order penalties induce smoother trends. In our study, we use k = 1, consistent with traditional volatility models like GARCH and SV, which typically assume stationary dynamics with abrupt clustering. Nevertheless, higher-order differencing may be better suited for applications with more gradually evolving volatility.
In practice, volatility processes may exhibit both local irregularities that are neither smooth nor abrupt. To account for this, we extend ASV by introducing a nugget effect, an additional noise term that absorbs small scale lodl fluctuations. We refer to this model as ASV with nugget:
$h_{t}=h^{*}_{t}+u_{t}$ $[u_{t}|\sigma^{2}_{c}]\sim N(0,\sigma^{2}_{c}),$ (5)
$[\Delta^{k}h_{t}^{*}|\sigma^{2}_{h,t}]\sim N(0,\sigma^{2}_{h,t})$ $[\log(\sigma^{2}_{h,t})|\mu,\phi]\sim DSP(a,b,\mu,\phi).$
ASV with nugget described in Equation 5 shares the same observation model in Equation 4, but differs in that the log-variance $h_{t}$ is now decomposed into a smooth component $h_{t}^{*}$ and a local noise term $u_{t}$, allowing the model to capture both structured and unstructured variation.
3 Bayesian Inference via Data Augmentation
We briefly discuss the Bayesian inference procedure for the ASV model described in Equation 4. The observed data are denoted by $\boldsymbol{y}:=(y_{1},\ldots,y_{T})^{\prime}$, the latent log-variance process by $\boldsymbol{h}:=(h_{1},\ldots,h_{T})^{\prime}$ and its associated log-variance process by $\boldsymbol{v}:=(\log(\sigma^{2}_{h,1}),\ldots,\log(\sigma^{2}_{h,T}))^{\prime}$. For the autoregressive parameter $\phi$, a shifted Beta prior: $(\phi+1)/2\sim Beta(1/2,1/2)$ is assumed. For the global location parameter, we assume $\mu\sim Z(1/2,1/2,0,1)$, which implies $\tau=\exp\{\mu(1-\phi)/2\}\sim C^{+}(0,1)$ when $\phi=0$. This recovers the global scale prior used in the horseshoe formulation by Carvalho et al. (2010). Inference proceeds by sampling from the joint posterior distribution of $(\boldsymbol{h},\boldsymbol{v},\mu,\phi)$ conditional on $\boldsymbol{y}$ via Gibbs sampling scheme that iteratively samples from the full conditional distributions.
Sampling from the conditional posterior distribution of $\boldsymbol{h}$ in the SV model is challenging due to its non-linear likelihood. While methods like sequential Monte Carlo by Jacquier et al. (1994) can sample $h_{t}$ from the exact conditional posterior sequentially, this approach is computationally demanding. Instead, we adopt a quasi-likelihood approach based on the transformed observations $\boldsymbol{y^{*}}:=(\log(y^{2}_{1}),\ldots,\log(y^{2}_{T}))^{\prime}$, recasting the model into a linear system with a non-Gaussian error term: $y_{t}^{*}=h_{t}+\log(\epsilon_{t}^{2}),\quad\epsilon_{t}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$.
The asymmetric error term $\log(\epsilon_{t}^{2})$ can be accurately approximated by a finite Gaussian mixture. The idea is originally proposed by Kim et al. (1998) and later refined by Omori et al. (2007), who introduced a 10-component Gaussian mixture approximation:
$\log(\epsilon_{t}^{2})\approx m_{j_{t}}+o_{t}$ $[o_{t}]\stackrel{{\scriptstyle iid}}{{\sim}}N(0,w_{j_{t}}^{2}),$
where $j_{t}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Categorical}(\pi^{\text{Omori}})$, with fixed weights $\pi^{\text{Omori}}=(\pi_{1},\ldots,\pi_{10})$. Each latent index $j_{t}\in\{1,\ldots,10\}$ selects a component-specific mean $m_{j_{t}}$ and standard deviation $w_{j_{t}}$ from a pre-specified table (Omori et al., 2007). This approximation transforms the quasi-likelihood into a conditionally Gaussian model, making posterior sampling significantly more efficient. Conditional on $j_{t}$, the full conditional distribution of $\boldsymbol{h}$ becomes a Gaussian likelihood with a structured Gaussian prior:
$y_{t}^{*}\approx h_{t}+m_{j_{t}}+o_{t}$ $[o_{t}|s_{j_{t}}]\stackrel{{\scriptstyle ind}}{{\sim}}N(0,w_{j_{t}}^{2}),$
$[\Delta^{k}h_{t}|v_{t}]\sim N(0,\exp(v_{t}))$ $[v_{t}|\mu,\phi]\sim DSP(a,b,\mu,\phi).$
The same Gaussian mixture approximation and data augmentation are applied to $\log(({\Delta^{k}h_{t}})^{2})$, which shares the same distribution as $\boldsymbol{y}^{*}$. A corresponding latent variable $s_{t}$ is introduced.
The evolution equation of $v_{t}$ in Equation 3 may be interpreted as the SV model with a heavy-tailed noise distribution $\eta_{t}\sim Z(a,b,0,1)$. We leverage the Pólya-Gamma scale mixture representation of the $Z$-distribution (Barndorff-Nielsen et al., 1982), which states:
$[\eta_{t}|\xi_{t}]\sim N(\xi_{t}^{-1}(a-b)/2,\xi_{t}^{-1})$ $[\xi_{t}]\sim PG(a+b,0).$
Combined with the approximation on $\log(\Delta^{k}h_{t}^{2})$, the log-variance evolution becomes conditionally Gaussian:
$\log((\Delta^{k}h_{t})^{2})\approx v_{t}+m_{s_{t}}+r_{t}$ $[r_{t}|w_{s_{t}}]\stackrel{{\scriptstyle ind}}{{\sim}}N(0,w_{s_{t}}^{2}),$
$v_{t}=\mu+\phi(v_{t-1}-\mu)+\eta_{t}$ $[\eta_{t}|\xi_{t}]\stackrel{{\scriptstyle ind}}{{\sim}}N(\xi_{t}^{-1}(a-b)/2,\xi_{t}^{-1}).$
We sample $\boldsymbol{v}$ using the efficient AWOL sampler proposed by Kastner and Frühwirth-Schnatter (2014). Sampling from the Pólya-Gamma distribution is performed by truncating an infinite sum as shown in (Polson et al., 2013). Similarly, $\xi_{\mu}\sim PG(1,0)$ is introduced to represent $\mu\sim Z(1/2,1/2,0,1)$ as $\mu|\xi_{\mu}\sim N(0,\xi_{\mu}^{-1})$.
In addition to the four existing parameters $(\boldsymbol{h},\boldsymbol{v},\mu,\phi)$, four additional parameters are introduced: the mixture indicators $\boldsymbol{j}:=(j_{1},\ldots,j_{T})^{\prime}$ and $\boldsymbol{s}:=(s_{1},\ldots,s_{T})^{\prime}$, and the Pólya-Gamma variables: $\boldsymbol{\xi}:=(\xi_{1}\ldots\xi_{T})^{\prime}$ and $\xi_{\mu}$. This section shows how the data augmentation framework enables efficient posterior sampling of $\boldsymbol{h}$ and $\boldsymbol{v}$. Detailed derivations of the full conditional distributions are discussed in the Appendix B.
For the nugget model described in Equation 5, the same approximation and data augmentation are applied. An additional Gaussian layer $h_{t}|h_{t}^{*},\sigma^{2}_{c}\sim N(h_{t}^{*},\sigma^{2}_{c})$ is introduced and the prior is placed on $\Delta^{k}h^{*}_{t}$ instead of $\Delta^{k}h_{t}$. However, the resulting model remains conditionally Gaussian, and the structure of the Gibbs sampler remains largely unchanged.
4 Simulation Study
4.1 Set-up
DGP1: SV with 1 Regime DGP2: SV with 2 Regimes
$y_{t}\sim N(0,\exp(h_{t}))$ $h_{t}=3+0.8(h_{t-1}-3)+0.2u_{t}$ $u_{t}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ $y_{t}\sim N(0,\exp(h_{t}))$ $h_{t}=m_{s_{t}}+0.8(h_{t-1}-m_{s_{t-1}})+0.2u_{t}$ $u_{t}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ $m_{s_{t}}=\begin{cases}\,-10\,,&\text{if $s_{t}=0$.}\\ \,6\,,&\text{if $s_{t}=1$.}\end{cases}$
DGP3: SV with 3 Regimes DGP4: GARCH with 1 Regime
$y_{t}\sim N(0,\exp(h_{t}))$ $h_{t}=m_{s_{t}}+0.8(h_{t-1}-m_{s_{t-1}})+0.2u_{t}$ $u_{t}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ $m_{s_{t}}=\begin{cases}\,-10\,,&\text{if $s_{t}=0$.}\\ \,-3\,,&\text{if $s_{t}=1$.}\\ \,3\,,&\text{if $s_{t}=2$.}\end{cases}$ $y_{t}\sim N(0,\sigma^{2}_{y,t})$ $\sigma_{y,t}^{2}=1+0.1y_{t-1}^{2}+0.5\sigma_{y,t-1}^{2}$
DGP5: GARCH with 2 Regimes DGP6: GARCH with 3 Regimes
$y_{t}\sim N(0,\sigma^{2}_{y,t})$ $\sigma_{y,t}^{2}=m_{s_{t}}+0.15y_{t-1}^{2}+\beta_{s_{t}}\sigma_{y,t-1}^{2}$ $m_{s_{t}}=\begin{cases}\,8\,,&\text{if $s_{t}=0$.}\\ \,0.1\,,&\text{if $s_{t}=1$.}\end{cases}$ $\beta_{s_{t}}=\begin{cases}\,0.8\,,&\text{if $s_{t}=0$.}\\ \,0.3\,,&\text{if $s_{t}=1$.}\end{cases}$ $y_{t}\sim N(0,\sigma^{2}_{y,t})$ $\sigma_{y,t}^{2}=m_{s_{t}}+0.15y_{t-1}^{2}+\beta_{s_{t}}\sigma_{y,t-1}^{2}$ $m_{s_{t}}=\begin{cases}\,12\,,&\text{if $s_{t}=0$.}\\ \,8\,,&\text{if $s_{t}=1$.}\\ \,0.1\,,&\text{if $s_{t}=2$.}\end{cases}$ $\beta_{s_{t}}=\begin{cases}\,0.8\,,&\text{if $s_{t}=0$.}\\ \,0.5\,,&\text{if $s_{t}=1$.}\\ \,0.2\,,&\text{if $s_{t}=2$.}\end{cases}$
DGP7: Sinusoidal DGP8: Piecewise Constant
$y_{t}\sim\mathcal{N}(0,\exp(h_{t}))$ $h=A\sin(10(2\pi t)/T)+B\cos(10(2\pi t)/T)$ $\quad+C\sin(3(2\pi t)/T)+D\cos(3(2\pi t)/T)$ $A,B,C,D\stackrel{{\scriptstyle ind}}{{\sim}}U(0,5)$ $y_{t}\sim\mathcal{N}(0,\exp(h_{t}))$ $j:=\left\lfloor\frac{t}{25}\right\rfloor+1$ $h_{t}=(-1)^{j}\cdot|z_{j}|$ $z_{j}\sim\begin{cases}N(5,0.5^{2})&\text{if $j$ is even}\\ N(0,0.5^{2})&\text{if $j$ is odd}\end{cases}$
Table 1: Summary of the eight data‐generating processes (DGP). DGP 1 - 3 simulate a stochastic‐volatility model with 1, 2, or 3 regimes; DGP 4–6 simulate a GARCH(1,1) model with 1, 2, or 3 regimes. In DGP 7, $h_{t}$ is a random sinusoid; in DGP 8, $h_{t}$ is piecewise constant with random intercepts. For DGPs 2 and 4, the transition probability of remaining in the same state is 0.98, and transitioning to a different state is 0.02. For DGPs 3 and 6, the probability of staying in the same state is 0.98, while switching to any other state occurs with probability 0.01. Each DGP comprises 1,000 sample paths, each path being 1,000 data points in length.
The simulation study evaluates the performance of the proposed ASV model against several existing methods, including the SV, Markov-Switching SV with two regimes (MSSV2), RWSV, GARCH, and two-regime GARCH (MSGARCH2). A local-only shrinkage model RWSV-BL is also included for comparison. Four variants of ASV: ASV with the horseshoe prior (ASV-HS), ASV with the dynamic horseshoe prior (ASV-DHS), and their respective versions with a nugget effect (ASV-HS-N and ASV-DHS-N) are considered. ASV-HS corresponds to the special case of Equation 4 with $a=b=1/2$ and $\phi=0$, while ASV-DHS generalizes this by allowing $\phi$ to be estimated from the data. The distinction between ASV with and without nugget is shown in Equation 5. All ASV models use $k=1$, applying shrinkage to the first-order differences of the latent log-variance process. This choice reflects the empirical observation that volatility in the SV or GARCH-type models typically exhibits clustering rather than gradual changes in slope.
Eight DGPs are considered: DGP 1–3 use SV models with one to three regimes; DGP 4–6 use GARCH models with one to three regimes; DGP 7 features sinusoidal log-variance with random coefficients; and DGP 8 involves piecewise constant log-variance with 3 to 5 random breakpoints. For each DGP, 1,000 sample paths are simulated. Detailed parameterization of the DGPs may be found in Table 1.
Model performance is evaluated using three metrics: Mean Absolute Error (MAE), Empirical Coverage (EC), and Mean Credible Interval Width (MCIW):
$MAE$ $=\frac{1}{T}\sum_{t=1}^{T}|\sigma_{t}-\hat{\sigma_{t}}|,$ $EC$ $=\frac{1}{T}\sum_{t=1}^{T}\mathds{1}_{(\hat{\sigma}_{t,0.05},\hat{\sigma}_{t,0.95})}(\sigma_{t}),$ $MCIW$ $=\frac{1}{T}\sum_{t=1}^{T}(\hat{\sigma}_{t,0.95}-\hat{\sigma}_{t,0.05}).$
The MAE measures the accuracy of the volatility estimate by calculating the average absolute difference between the true volatility $\sigma_{t}$ and the model’s estimate $\hat{\sigma_{t}}$. For Frequentist methods, the point estimate is based on its maximum likelihood estimate, while it is the posterior mean for the Bayesian methods. The EC assesses the reliability of each model’s uncertainty quantification by calculating the proportion of the true volatility falling within the model’s 90% credible interval. Finally, the MCIW evaluates the precision of the uncertainty quantification, where a narrower credible interval represents a higher level of precision.
All models are implemented in R language (R Core Team, 2013). Specifically, the SV model is implemented via stochvol package (Kastner, 2016), MSSV2 as specified in Hamilton (1989) and RWSV are implemented by the authors as no readily available packages in R exist. GARCH(1,1) is fitted via fGarch package (Wuertz et al., 2023), and MSGARCH2 is fitted via MSGARCH package (Ardia et al., 2019). For fitting RWSV-BL, dsp package by Kowal et al. (2019) and genlasso package by Arnold and Tibshirani (2022) are used. dsp package by Kowal et al. (2019) is expanded by the authors to implement ASV. For Bayesian models, 5,000 posterior samples are generated after 20,000 burn-ins for each sample path.
4.2 Results
DGP ID
1 2 3 4 5 6 7 8
True Model SV (1 regime) SV (2 regimes) SV (3 regimes) GARCH (1 regime) GARCH (2 regimes) GARCH (3 regimes) Sinusoidal Piecewise
MAE SV 0.4183 (0.0195) 4.3355 (0.9853) 0.4948 (0.1645) 0.1641 (0.0212) 1.2965 (0.2978) 1.2804 (0.3159) 1.1831 (0.9398) 2.0796 (0.1713)
MSSV2 0.6288 (1.6236) 1.0265 (0.282) 0.4904 (0.1676) 0.3034 (1.1001) 2.6922 (2.9965) 3.1646 (3.4242) 2.6815 (2.5832) 2.019 (0.1632)
GARCH 0.4297 (0.0227) 5.8768 (0.817) 0.9623 (0.2776) 0.0421 (0.0181) 2.0226 (0.456) 1.8237 (0.437) 2.7572 (2.2338) 4.1509 (0.3525)
MSGARCH2 0.4334 (0.0318) 5.9983 (1.6226) 1.1078 (1.2224) 0.0577 (0.0206) 0.9559 (0.1991) 1.247 (0.26) 2.1427 (1.6771) 2.7224 (0.7748)
RWSV 0.4745 (0.0309) 4.2692 (0.9614) 0.4875 (0.161) 0.1704 (0.0199) 1.2744 (0.2905) 1.2576 (0.309) 1.1668 (0.925) 2.0301 (0.1692)
RWSV-BL 1.0103 (0.0597) 3.4705 (0.7682) 0.4641 (0.152) 0.2811 (0.0311) 1.4004 (0.3371) 1.4543 (0.3761) 1.42 (1.0283) 1.9933 (0.1658)
[…]
ASV-DHS 0.4517 (0.0313) 1.4892 (0.3192) 0.2369 (0.0725) 0.1777 (0.0212) 0.9462 (0.2124) 1.0336 (0.2535) 0.9426 (0.7442) 1.2209 (0.1385)
ASV-HS-N 0.4437 (0.0292) 1.4623 (0.3275) 0.236 (0.0746) 0.1776 (0.0209) 0.9754 (0.2228) 1.0617 (0.2588) 0.9358 (0.7373) 1.2449 (0.1386)
ASV-DHS-N 0.4477 (0.0304) 1.4892 (0.3205) 0.237 (0.0725) 0.1783 (0.0212) 0.9537 (0.2146) 1.0414 (0.2558) 0.9427 (0.7423) 1.2242 (0.1382)
MCIW SV 1.7151 (0.4996) 23.9914 (5.5341) 2.6681 (0.943) 0.8261 (0.0985) 6.9493 (1.6761) 6.9645 (1.7686) 6.9224 (5.4048) 10.5937 (0.8248)
MSSV2 1.6856 (0.8774) 6.1165 (2.2849) 2.9168 (0.9828) 1.2632 (1.4566) 8.1412 (1.9785) 8.4577 (2.134) 7.9837 (4.4592) 11.0375 (0.8436)
GARCH - - - - - - - -
MSGARCH2 - - - - - - - -
RWSV 1.4682 (0.0678) 23.6249 (5.4109) 2.6127 (0.9215) 0.707 (0.1036) 6.7469 (1.6199) 6.7313 (1.7063) 6.801 (5.3228) 10.0901 (0.7914)
RWSV-BL 5.8851 (0.178) 19.7094 (4.3992) 2.5829 (0.8688) 1.6936 (0.1421) 7.9766 (1.9155) 8.479 (2.1574) 8.4838 (6.0041) 10.6267 (0.7472)
[…]
ASV-DHS 1.0711 (0.1338) 6.5254 (1.3109) 1.026 (0.3029) 0.6442 (0.1047) 4.0214 (0.8943) 3.9701 (0.9812) 4.3191 (3.3025) 6.3782 (0.4163)
ASV-HS-N 1.7522 (0.3342) 7.0255 (1.4975) 1.1211 (0.354) 0.7136 (0.1034) 4.4482 (1.0088) 4.4818 (1.0818) 4.6034 (3.4542) 6.6306 (0.4301)
ASV-DHS-N 1.7898 (0.3295) 7.1332 (1.431) 1.1215 (0.3332) 0.7127 (0.1017) 4.2736 (0.9445) 4.3395 (1.0468) 4.6343 (3.4943) 6.5572 (0.4303)
EC SV 0.8655 (0.0968) 0.9159 (0.0093) 0.9177 (0.0104) 0.9621 (0.0131) 0.9296 (0.0116) 0.9327 (0.0121) 0.9764 (0.0114) 0.9091 (0.0107)
MSSV2 0.8348 (0.1264) 0.9589 (0.0368) 0.939 (0.0113) 0.9839 (0.0176) 0.9451 (0.0105) 0.9514 (0.0103) 0.9757 (0.0085) 0.9173 (0.0101)
GARCH - - - - - - - -
MSGARCH2 - - - - - - - -
RWSV 0.7779 (0.0276) 0.9144 (0.0094) 0.9148 (0.0107) 0.9081 (0.0248) 0.9252 (0.0123) 0.9274 (0.013) 0.9758 (0.0118) 0.8973 (0.0112)
RWSV-BL 0.9718 (0.0098) 0.9549 (0.0092) 0.9515 (0.0101) 0.972 (0.01) 0.9557 (0.0099) 0.96 (0.0097) 0.975 (0.0093) 0.9422 (0.0107)
ASV-HS 0.6221 (0.0435) 0.9106 (0.0265) 0.9094 (0.0263) 0.8678 (0.0468) 0.9194 (0.0456) 0.8721 (0.0582) 0.9275 (0.0295) 0.9569 (0.0168)
ASV-DHS 0.6426 (0.041) 0.9009 (0.0281) 0.8973 (0.0278) 0.8584 (0.045) 0.9186 (0.0387) 0.8765 (0.0519) 0.9264 (0.0296) 0.9583 (0.017)
ASV-HS-N 0.8701 (0.0577) 0.9388 (0.0206) 0.9371 (0.0205) 0.9014 (0.0335) 0.9317 (0.0374) 0.8996 (0.0448) 0.9498 (0.0245) 0.9614 (0.0156)
ASV-DHS-N 0.8753 (0.0539) 0.9355 (0.0217) 0.9329 (0.0219) 0.8988 (0.0314) 0.9338 (0.032) 0.9035 (0.0415) 0.9494 (0.0245) 0.963 (0.0155)
Table 2: Model performance under each data-generating process (DGP 1–8), evaluated using: MAE (mean absolute error), MCIW (mean 90% credible interval width), and EC (empirical 90% coverage). Standard deviations are shown in parentheses. For each DGP, the two models with the lowest average MAE and the four models with the narrowest average MCIW are highlighted in bold.
(a) DGP 1
(b) DGP 2
(c) DGP 3
(d) DGP 4
(e) DGP 5
(f) DGP 6
(g) DGP 7
(h) DGP 8
Figure 3: Box plots of Mean Absolute Error (MAE) across 1,000 sample paths comparing Stochastic Volatility (SV), Markov-Switching Stochastic Volatility with 2 Regimes (MSSV2), Generalized AutoRegressive Conditional Heteroskedasticity (GARCH), Markov-Switching GARCH with 2 Regimes (MSGARCH2), Random Walk SV with Inverse Gamma Prior (RWSV), Random Walk SV with Bayesian LASSO (RWSV-BL), Adaptive SV with the horseshoe prior (ASV-HS), and ASV with dynamic horseshoe prior (ASV-DHS). Perfectly specified models, drawn in black, are SV and MSSV2 for DGP 1 and 2 and GARCH and MSGARCH2 for DGP 4 and 5. All models are misspecified for DGP 3 and 6.
As shown in Table 4.2 and Figure 3, the proposed ASV models consistently outperform or match the true models in DGPs with large volatility shifts, including DGPs 2, 3, 5, 6, and 8. In DGP 2, their MAE is only marginally higher than that of MSSV2, and in DGP 5, even slightly lower than that of MSGARCH2, both of which are correctly specified for those settings. Notable differences are found in DGPs 3, 6, and 8. In DGP 3, the average MAE of ASV models is 0.24 (sd = 0.07), compared to 0.46 (sd = 0.15) for the next best model, RWSV-BL, leading to 44% reduction. Similar differences are observed in DGPs 5 and 6. Such results underscore ASV’s ability in capturing abrupt structural changes.
ASV models also perform reasonably well in scenarios where the underlying volatility is stationary or smoothly varying, as in DGPs 1, 4, and 7. SV and GARCH achieve the lowest average MAEs in DGPs 1 and 4 due to perfect specification; ASV models perform similarly to other misspecified models. In DGP 7, where volatility evolves smoothly without abrupt shifts, ASV outperforms all models including SV and RWSV. This suggests that although ASV is designed to capture abrupt structural changes, it still maintains competitive performance in settings that favor smooth or mean-reverting dynamics.
Also, the simulation study highlights the limitations of HMS models, MSSV2 and MSGARCH2, which rely on a fixed number of latent regimes. These models perform poorly when the assumed regime structure does not align with the data. Notably, they perform worse than their single-regime counterparts SV and GARCH in scenarios with stationary volatility (DGPs 1 and 4). Their performance also deteriorates when the number discrete levels are more than 2 (DGPs 3, 6, and 8). Even in DGP 7, where volatility evolves smoothly without abrupt changes, the discrete regime-switching assumption introduces artificial segmentation, leading to subpar performance. In contrast, ASV models adaptively estimate volatility using global-local shrinkage, flexibly adjusting to a wide range of structural patterns without assuming a fixed number of regimes.
The second and third row of Table 4.2 presents the ECs and the MCIWs of the 90% credible intervals of the eight Bayesian methods in the simulation study: SV, MSSV2, RWSV, RWSV-BL and 4 variants of ASV. ASV-HS and ASV-DHS achieve nominal 90% coverage in all DGPs except in DGPs 1 and 4. In DGP 1, they exhibit substantial undercoverage, with empirical coverage dropping to 63%, and in DGP 4, mild undercoverage at 86%. This is likely due to the models’ assumption of nonstationarity, which conflicts with the stationary assumption of the DGP. ASV-HS-N and ASV-DHS-N address the undercoverage problem by introducing a nugget effect, achieving correct nominal coverage across all 8 DGPs. In addition, their credible intervals are the narrowest across all eight DGPs aside from their non-nugget counter parts, indicating both precise and robust uncertainty quantification.
The key structural difference between ASV-DHS and ASV-HS is the presence of an autoregressive shrinkage parameter $\phi$ in ASV-DHS, which governs the temporal evolution of the local scale parameters. Despite this, they exhibit similar performance across DGPs 1 through 7. This is expected: in DGPs 1, 4, and 7, changes in volatility are gradual. In DGPs 2, 3, 5, and 6, abrupt changes occur but are highly persistent with the transition probability between different states of 2% for DGPs 2 and 5, and only 1% for DGPs 3 and 6, resulting in relatively few shifts. In DGP 8 features frequent and periodic volatility changes. In this setting, ASV-DHS consistently achieves smaller MAEs than ASV-HS. Although the average improvement is modest (0.021), it is consistent across 1,000 replicates, indicating that dynamic shrinkage offers better local adaptivity when volatility shifts are both abrupt and recurrent.
5 Empirical Study
(a) S&P 500: SV
(b) S&P 500: MSSV2
(c) S&P 500: ASV-DHS
(d) S&P 500: ASV-DHS-N
(e) Electricity: SV
(f) Electricity: MSSV2
(g) Electricity: ASV-DHS
(h) Electricity: ASV-DHS-N
(i) Bike Rental: SV
(j) Bike Rental: MSSV2
(k) Bike Rental: ASV-DHS
(l) Bike Rental: ASV-DHS-N
Figure 4: Estimated log-variance from four methods, Stochastic Volatility (SV), Markov-Switching SV with two regimes (MSSV2), Adaptive Stochastic Volatility with the dynamic horseshoe prior (ASV-DHS), and ASV-DHS with a nugget effect, applied to three datasets: weekly S&P 500 log returns, daily electricity demand, and daily bike rental counts.
We continue our analysis of the datasets introduced in Section 1 and estimate the log-variance using four methods: SV and MSSV2, which serve as established baselines, and ASV-DHS and ASV-DHS-N, our proposed models. For the ASV models, we set $k=1$ to reflect the presence of abrupt changes observed in Figure 1. In the S&P 500 dataset (Figures 4(a), 4(b), 4(c) and 4(d)), all models detect major volatility spikes in 2009 and 2020, corresponding to the global financial crisis and the COVID-19 pandemic. Moderate increases are also identified around 2001 (9/11), 2011 (European debt crisis), and 2018 (U.S. government shutdowns). ASV-DHS and ASV-DHS-N provide smoother and more stable estimates compared to SV and MSSV2, particularly during low-volatility periods.
For both the electricity (Figures 4(e), 4(f), 4(g) and 4(h)), and bike rental datasets (Figures 4(i), 4(j), 4(k) and 4(l)), all methods capture strong seasonal patterns in volatility, with elevated uncertainty coinciding with periods of peak activity. In the electricity dataset, this corresponds to high-demand months, while in the bike rental dataset, volatility increases during summer.
Dataset Metric SV MSSV2 ASV-DHS ASV-DHS-N
S&P 500 $\mathrm{Mean}$ 0.0739 0.0792 0.037 0.0799
$\mathrm{Kurtosis}$ 0.9105 0.7348 52.3562 11.5863
$\mathrm{CP}$ 0 0 9 3
Electricity $\mathrm{Mean}$ 0.0307 0.0600 0.0176 0.0213
$\mathrm{Kurtosis}$ 0.1691 7.5788 65.4116 68.4159
$\mathrm{CP}$ 0 4 16 20
Bike $\mathrm{Mean}$ 0.0616 0.143 0.0456 0.1331
$\mathrm{Kurtosis}$ 0.1989 13.3789 90.7971 5.6332
$\mathrm{CP}$ 0 3 10 2
Table 3: Summary statistics for each dataset and model: (1) the mean and (2) kurtosis of the estimated log-variance increments $|\Delta\hat{h}_{t}|$ and (3) the total number of flagged change points (CP). CP are computed as $\sum_{t=2}^{T}\mathds{1}\{|\Delta\hat{h}_{t}|>\delta\}$, where $\delta:=5\times\texttt{sd}(\Delta\hat{h}_{t})$.
Table 3 presents summary metrics of the estimated log-variance increments $|\Delta\hat{h}_{t}|$ across the three empirical datasets. These metrics, mean, kurtosis, and the number of flagged change points (CP), offer insight into the smoothness, tail behavior, and local adaptivity. ASV-DHS consistently produces the lowest mean, indicating highly smooth volatility trajectories. At the same time, it shows extremely high kurtosis and the highest number of flagged change points, suggesting heavy-tailed behavior in the increments and strong local adaptivity to sudden structural shifts. These properties highlight the model’s tendency to promote overall smoothness while remaining responsive to large isolated changes.
The smoothness of ASV-DHS-N adapts to the structure of the data. In the electricity dataset, where the volatility is characterized by extended periods of small changes punctuated by occasional abrupt shifts, the mean in Table 3 under ASV-DHS-N is similar to that of ASV-HS. In contrast, for the bike rental and S&P 500 datasets, which exhibit more unpredictable changes in volatility, the smoothness profile of ASV-DHS-N resembles that of MSSV2. The kurtosis remains substantially higher than that of SV and MSSV2, preserving the model’s ability to capture heavy-tailed behavior. The CP is also notably higher than that of SV, reflecting greater local adaptivity, though still lower than that of ASV-DHS.
(a) S&P 500 Index
(b) Electricity
(c) Bike Rentals
Figure 5: Comparison of the posterior means of the shrinkage parameter $\kappa_{t}:=1/(1+\exp\{v_{t}\})$, where $v_{t}=\log(\mathrm{Var}(\Delta h_{t}))$, and the log-variance parameter $h_{t}$, estimated using ASV-DHS-N on the S&P 500, electricity demand, and bike rental datasets. The shaded regions represent the 95% and centered 90% credible intervals for $\kappa_{t}$ and $h_{t}$, respectively. Vertical lines indicate dates where the expected $\kappa_{t}$ falls below 0.9.
Another defining feature of ASV is its ability to localize the timing of large variations in volatility through the shrinkage parameter $\kappa_{t}:=\frac{1}{1+\exp(v_{t})}$. $\kappa_{t}$ close to 0 corresponds to high variability in the latent volatility increments $\Delta h_{t}$, indicating periods of substantial changes in $h_{t}$, while a $\kappa_{t}$ near 1 suggests strong shrinkage and minimal change. Thus $\kappa_{t}$ serves as an interpretable signal for identifying the timing of structural shifts in volatility.
In Figure 5, we highlight time points where the posterior mean of $\kappa_{t}$ drops below 0.9, marking potential locations of abrupt volatility changes. For the S&P 500 index (Figure 5(a)), this includes notable episodes explored in earlier paragraph, such as the 2008 financial crisis, the 2011 European debt crisis, and the COVID-19 shock in early 2020. In Figure 5(b), electricity usage spikes in volatility around late spring and early fall, periods that typically mark transitions in heating or cooling demand, reflecting changing usage patterns driven by weather fluctuations. Figure 5(b) shows a similar pattern with volatility spikes in bike rental counts occurring near the start of spring such as late February and March and near the end-of-year holiday season in December and January.
6 Trend Filtering Jointly in Mean and Variance
The Bayesian Trend Filter with DSP (BTF-DSP) by Kowal et al. (2019) provides a smooth and locally adaptive estimate of the mean of a time series. As an extension, we propose the Bayesian Trend Filter with ASV (BTF-ASV), thereby allowing locally adaptive estimates of both the mean $\{\beta_{t}\}_{t=1}^{T}$, and the log variance process $\{h_{t}\}_{t=1}^{T}$ simultaneously:
$y_{t}=\beta_{t}+\exp\{h_{t}/2\}\epsilon_{t}$ $[\epsilon_{t}]\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1),$ (6)
$[\Delta^{k_{\beta}}\beta_{t}|\sigma^{2}_{\beta,t}]\sim N(0,\sigma^{2}_{\beta,t})$ $[\log(\sigma^{2}_{\beta,t})|\mu_{\beta},\phi_{\beta}]\sim DSP(a_{\beta},b_{\beta},\mu_{\beta},\phi_{\beta}),$
$[\Delta^{k_{h}}h_{t}|\sigma^{2}_{h,t}]\sim N(0,\sigma^{2}_{h,t})$ $[\log(\sigma^{2}_{h,t})|\mu_{h},\phi_{h}]\sim DSP(a_{h},b_{h},\mu_{h},\phi_{h}).$
We impose two conditionally independent priors on the kth-order differences of $\beta_{t}$ and $h_{t}$. The model distinguishes changes in the mean from changes in volatility through its structural formulation: $\beta_{t}$ enters the likelihood additively, while $h_{t}$ governs the multiplicative noise scale. This separation ensures that fluctuations in the variance do not influence the location of the distribution. Below, we show that the model is identifiable in the sense that the evolution variance terms, $\sigma^{2}_{\beta,t}$ and $\sigma^{2}_{h,t}$, are uniquely determined by the observed data.
Theorem 4.
The following model is identifiable:
$y_{t}=\beta_{t}+\exp\{h_{t}/2\}\epsilon_{t}$ $\quad[\epsilon_{t}]\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1),$
$[\Delta^{k}\beta_{t}|\sigma^{2}_{\mu,t}]\sim N(0,\sigma_{\mu,t}^{2})$ $\quad[\Delta^{k}h_{t}|\sigma^{2}_{h,t}]\sim N(0,\sigma_{h,t}^{2}).$
Proof.
For contradiction, fix two distinct parameter vectors $\boldsymbol{\theta}:=(\boldsymbol{\beta},\boldsymbol{h},\boldsymbol{\lambda}_{\mu},\boldsymbol{\lambda}_{h})$ and $\boldsymbol{\theta}^{*}:=(\boldsymbol{\beta}^{*},\boldsymbol{h}^{*},\boldsymbol{\lambda}_{\mu}^{*},\boldsymbol{\lambda}_{h}^{*})$ such that $\boldsymbol{\theta}\neq\boldsymbol{\theta}^{*}$ and $f(y|\boldsymbol{\theta})=f(y|\boldsymbol{\theta}^{*})$. Indeed, $\mathrm{Var}(y|\boldsymbol{\theta})=\mathrm{Var}(y|\boldsymbol{\theta}^{*}).$ This covariance equality necessarily forces $\boldsymbol{\theta}=\boldsymbol{\theta}^{*}$, leading to contradiction. Detailed proof is shown in Appendix A. ∎
(a) Observed Series $y_{t}$
(b) Estimated $\beta_{t}$ with 90% CR
(c) Estimated $h_{t}$ with 90% CR
Figure 6: Monthly mean total sunspot number, the average value of the daily total sunspot numbers over all days of each calendar month, from Jan. 1901 to Aug. 2024 is illustrated in 6(a). The Bayesian Trend Filter with Adaptive Stochastic Volatility (BTF-ASV) is used to estimate the mean trend $\beta_{t}$ and is illustrated in 6(b). The estimate of log-variance $h_{t}$ and its 90% credible intervals is shown in 6(c).
BTF-ASV is applied to the sunspot data to estimate both the time varying mean and the volatility. The data consist of monthly mean total sunspot numbers, defined as the average daily total over each calendar month, from January 1901 to August 2024, obtained from SILSO World Data Center (2025). The number of sunspots exhibits a well-known cyclic pattern associated with the 11-year solar cycle, which reflects periodic fluctuations in solar magnetic activity. This cyclical behavior is clearly visible in the raw observations shown in Figure 6(a). The BTF estimate of the mean trend, $\beta_{t}$, shown in Figure 6(b), provides a smooth reconstruction of this underlying cycle.
Interestingly, the estimated log-variance (Figure 6(c)) also follows a cyclical pattern that closely aligns with the mean trend. Periods of increased solar activity are associated with elevated volatility, suggesting that not only does solar activity rise and fall periodically, but the variability around these phases also intensifies. This highlights BTF-ASV’s ability to decompose a noisy data into structured patterns in both the mean and variance process, offering a richer understanding of solar dynamics.
7 Conclusion
We propose ASV, a flexible and locally adaptive model for capturing both abrupt and gradual changes in volatility. ASV extends RWSV by incorporating global-local shrinkage priors, specifically DSP by Kowal et al. (2019). We develop and study multiple variants of ASV, including models with the horseshoe (ASV-HS), dynamic horseshoe (ASV-DHS), and their nugget-augmented counterparts (ASV-HS-N and ASV-DHS-N).
Through extensive simulation studies, we show that ASV-HS-N and ASV-DHS-N consistently achieve low MAE, correct nominal coverage, and narrower credible intervals compared to existing volatility models. Although ASV variants without the nugget term perform similarly in terms of point estimation, they tend to exhibit under-coverage in stationary settings due to their lack of a noise-absorbing layer. Empirical applications demonstrate that ASV models effectively capture large structural changes in volatility while preserving smooth estimates in stable periods. Moreover, the shrinkage parameter $\kappa_{t}$ provides a useful diagnostic for further identifying the timing of structural shifts.
Finally, we integrate ASV into the error process of the Bayesian Trend Filter proposed in Kowal et al. (2019), resulting in BTF-ASV. Empirical studies reveal its ability to produce smooth estimates of both time-varying mean and time-varying variance. Future research directions could explore the applications of ASV and BTF-ASV in diverse fields such as finance, environmental science, and epidemiology, leveraging their adaptability in modeling time-series exhibiting complex patterns.
8 Acknowledgement
Financial support from National Science Foundation grants OAC-1940124 and DMS-2114143 is gratefully acknowledged.
9 Supplementary Material
Appendices:
A document including detailed proofs on the properties of Dynamic Shrinkage Processes (DSP) in section A. The exact prior distributions and detailed derivations of the full conditional distributions for Gibbs sampling in section B.
Appendix A Proofs of Theorems
A.1 Theorem 1
As mentioned, proofs directly follow from Kolmogorov’s three series theorem, which state the sufficient and necessary condition for almost sure convergence of infinite sum of random variables. Let $\phi\in(0,1)$. Due to symmetry of $sech$ function, we only need to prove the case when $10<\phi<$ 1, and the case where $-1<\phi<0$, directly follows. For the first condition, let $\epsilon>0$, and $h$ is some positive integer or 0.
[…]
$=\frac{4}{\pi}\text{arctan}(\text{exp}\{x/(2\phi^{h})\})\bigg{]}_{\epsilon}^{\infty}$
$=2(1-\frac{2}{\pi}\text{arctan}(\text{exp}\{\epsilon/(2\phi^{h})\})$
$=\frac{4}{\pi}\text{arccotan}(\text{exp}\{\epsilon/(2\phi^{h})\}$
$\leq\frac{4}{\pi}\text{arccotan}(\text{exp}\{h\})$ if $\textit{if }\epsilon\leq 2h\phi^{h}$
Since $10<\phi<$ 1, $2h\phi^{h}$ converges to 0. Also, $2h\phi^{h}\geq 0$ for $h\geq 0$. For any $\epsilon$, we have $h^{\prime}$ that satisfies $\epsilon\leq 2h\phi^{h}$ for $h^{\prime}\leq h$. Then, fix such $h^{\prime}$.
[…]
$\sum_{h=h^{\prime}}^{\infty}\frac{4}{\pi}\text{arccotan}(\text{exp}\{\epsilon/(2\phi^{h})\}\leq\sum_{h=h^{\prime}}^{\infty}\frac{4}{\pi}\text{arccotan}(\text{exp}\{h\})<\infty$
Therefore,
$\sum_{h=0}^{\infty}\mathrm{P}(|z_{h}|\geq\epsilon)=\sum_{h=0}^{\infty}\frac{4}{\pi}\text{arccotan}(\text{exp}\{\epsilon/(2\phi^{h})\}<\infty$
On the side note, we can also easily see that when $|\phi|>1$, the sum diverges to $\infty$. By Borel-Cantelli lemma, we may conclude that $\sum_{h=0}^{\infty}z_{h}$ diverges almost surely.
Define $y_{h}=z_{h}1_{\{|z_{h}|\leq\epsilon\}}$. For the second and the third condition, we need to show that 1) $\sum_{h=0}^{\infty}\mathbb{E}(y_{h})$ converges, and 2) $\sum_{h=0}^{\infty}\mathrm{Var}(y_{h})$, converges. For the second condition, it suffices to show: E(∑_h=0^∞ y_h) = ∑_h=0^∞ E(y_h) ¡ ∞, which is satisfied by showing $\sum_{h=0}^{\infty}\mathbb{E}(|y_{h}|)<\infty$ (Fubini). Using Cauchy-Schwartz Inequality:
$\sum_{h=0}^{\infty}\mathbb{E}(|y_{h}|)$ $=\sum_{h=0}^{\infty}\mathbb{E}(|z_{h}1_{\{|z_{h}|\leq\epsilon\}}|)$
$\leq\sum_{h=0}^{\infty}\sqrt{\mathbb{E}(z_{h}^{2})\mathrm{P}(|z_{h}|\leq\epsilon)}$
$\leq\sum_{h=0}^{\infty}\sqrt{\mathbb{E}(z_{h}^{2})}=\sum_{h=0}^{\infty}\sqrt{\mathrm{Var}(z_{h})}=\pi\sum_{h=0}^{\infty}\phi^{h}=\frac{\pi}{(1-\phi)}<\infty.$
The second condition is satisfied. For the third condition, similar logic is applied. First note that:
$\mathrm{Var}(\sum_{h=0}^{\infty}y_{h})$ $=\mathbb{E}((\sum_{h=0}^{\infty}y_{h})^{2})+\mathbb{E}((\sum_{h=0}^{\infty}y_{h}))^{2}$
$=\mathbb{E}((\sum_{h=0}^{\infty}y_{h})^{2})+((\sum_{h=0}^{\infty}\mathbb{E}(y_{h}))^{2}$
$=\mathbb{E}((\sum_{h=0}^{\infty}y_{h})^{2}).$
And,
$\lim_{h\to\infty}\mathbb{E}((\sum_{n=0}^{n}y_{h})^{2})$ $=\lim_{n\to\infty}\mathbb{E}(\sum_{h=0}^{n}y_{h}^{2})$ $(independence)$
$=\lim_{n\to\infty}\sum_{h=0}^{n}\mathbb{E}(y_{h}^{2})=\sum_{h=0}^{\infty}\mathbb{E}(z_{h}^{2}1_{\{|z_{h}|\leq\epsilon\}})$
$\leq\sum_{h=0}^{\infty}\mathbb{E}(z_{h}^{2})=\frac{\pi^{2}}{(1-\phi^{2})}<\infty$
Therefore:
$\mathrm{Var}(\sum_{h=0}^{\infty}y_{h})$ $=\sum_{h=0}^{\infty}\mathrm{Var}(y_{h})<\infty$
All three conditions are satisfied.
A.2 Theorem 2
When $\phi=0.5$, we have the following special case for the MGF,
$\prod_{h=0}^{\infty}\cos(\frac{\pi t}{2^{h}})$ $=\prod_{h=0}^{\infty}\frac{\sin(\frac{\pi t}{2^{h-1}})}{2\sin(\frac{\pi t}{2^{h}})}=\lim_{n\to\infty}\prod_{h=0}^{n}\frac{\sin(\frac{\pi t}{2^{h-1}})}{2\sin(\frac{\pi t}{2^{h}})}$
$=\lim_{n\to\infty}\frac{\sin(2\pi t)}{2^{n}\sin(\frac{\pi t}{2^{n}})}=\frac{\sin(2\pi t)}{2\pi t}$
We have:
$\prod_{h=0}^{\infty}\sec(\frac{\pi t}{2^{h}})=\frac{1}{\prod_{h=0}^{\infty}\cos(\frac{\pi t}{2^{h}})}=\frac{2\pi t}{\sin(2\pi t)}=\Gamma(1-2t)\Gamma(1+2t)=\mathrm{B}(1-2t,1+2t)$
$\frac{2\pi t}{\sin(2\pi t)}=\Gamma(1-2t)\Gamma(1+2t)$ is by the reflection relation. $Logistic(\mu,s)$ has the the following moment generating function $\text{exp}(-\mu t)\mathrm{B}(1-st,1+st)$. Since the moment generating function uniquely determines the random variable, and by Theorem 2,
$z_{t}=\sum_{h=0}^{\infty}z_{t,h}\stackrel{{\scriptstyle a.s}}{{\rightarrow}}Logistic(0,2)$
A.3 Theorem 3
Define $\lambda_{t}:=\text{exp}(v_{t})$. We showed that $f(\lambda_{t})=\frac{1}{(1+\lambda_{t})^{2}}$. We have:
[…]
$\forall x>0$, 12(1+x2)1(1+x)21(1+x2). Thus,
[…]
$=\frac{1}{2\sqrt{2\pi}}\text{exp}(\Delta^{2}h_{t}/2)E_{1}(\Delta^{2}h_{t}/2),$
where $E_{1}()$ is the exponential integral function, which satisfies the following upper and lower bound $\forall t>0$: exp(-t)2 log(1 + 2t) ¡ E_1(t) ¡ exp(-t) log(1 + 1t) Thus, f(Δh_t) ¡ 122 π log(1 + 2Δ2ht) Similarly for the lower bound, f(Δh_t) ¿ 182 π log(1 + 4Δ2ht) The lower and upper bound for $\Delta h_{t}\neq 0$ is shown. The lower bound clearly approaches infinity as $\Delta h_{t}$ approaches 0, which completes the proof.
A.4 Theorem 4
To prove identifiability of the hierarchical model, we need to establish that
  1. 1.
    For any $(\boldsymbol{h},\boldsymbol{\beta})$ and $(\boldsymbol{h}^{*},\boldsymbol{\beta}^{*})$ in the parameter space, if $f(\boldsymbol{y}|\boldsymbol{h},\boldsymbol{\beta})=f(\boldsymbol{y}|\boldsymbol{h}^{*},\boldsymbol{\beta}^{*}),\forall\boldsymbol{y}\in\mathcal{Y}$, then $(\boldsymbol{h},\boldsymbol{\beta})=(\boldsymbol{h}^{*},\boldsymbol{\beta}^{*}).$
  2. 2.
    For any $(\boldsymbol{\lambda_{\beta}},\boldsymbol{\lambda_{h}})$ and $(\boldsymbol{\lambda_{\beta}}^{*},\boldsymbol{\lambda_{h}}^{*})$ in the hyperparameter space, if $f(\boldsymbol{y}|\boldsymbol{\lambda_{\beta}},\boldsymbol{\lambda_{h}})=f(\boldsymbol{y}|\boldsymbol{\lambda_{\beta}}^{*},\boldsymbol{\lambda_{h}}^{*}),\forall\boldsymbol{y}\in\mathcal{Y}$, then $(\boldsymbol{\lambda_{\beta}},\boldsymbol{\lambda_{h}})=(\boldsymbol{\lambda_{\beta}}^{*},\boldsymbol{\lambda_{h}}^{*}).$
The first statement follows from the fact that, conditional on $(\boldsymbol{\beta},\boldsymbol{h})$, the likelihood is Multivariate Gaussian with mean $\boldsymbol{\beta}$ and with diagonal covariance matrix $\text{exp}\{\boldsymbol{h}\}$. Two parameters affect the density in functionally independent ways, thus the statement follows.
To prove the identifiability of the hyperparameter pair $(\boldsymbol{\lambda}_{\beta},\boldsymbol{\lambda}_{h})$, we proceed by contradiction. Suppose there exist two distinct parameter vectors $(\boldsymbol{\lambda}_{\beta},\boldsymbol{\lambda}_{h})\neq(\boldsymbol{\lambda}_{\beta}^{*},\boldsymbol{\lambda}_{h}^{*})$ such that $f(y|\boldsymbol{\lambda}_{\beta},\boldsymbol{\lambda}_{h})=f(y|\boldsymbol{\lambda}_{\beta}^{*},\boldsymbol{\lambda}_{h}^{*})$.
By model specification $\boldsymbol{\beta}|\boldsymbol{\lambda}_{\beta}\sim N(\boldsymbol{0},\mathbf{S}\Lambda_{\beta}\mathbf{S}^{T})$ and $\boldsymbol{h}|\boldsymbol{\lambda}_{h}\sim N(\boldsymbol{0},\mathbf{S}\Lambda_{h}\mathbf{S}^{T})$ where $\Lambda_{\beta}=diag(\boldsymbol{\lambda}^{2}_{\beta})$, $\Lambda_{h}=diag(\boldsymbol{\lambda}^{2}_{h})$ and $\mathbf{S}$ is the un-differencing operator matrix corresponding to the differencing order k in $\Delta^{k}\beta$ and $\Delta^{k}h$.
Using the law of total variance, the variance of $y$ conditional on the hyperparameters can be decomposed as:
$\mathrm{Var}(y\mid\boldsymbol{\lambda}_{\beta},\boldsymbol{\lambda}_{h})$ $=E_{\boldsymbol{\beta},\boldsymbol{h}\mid\boldsymbol{\lambda}_{\beta},\boldsymbol{\lambda}_{h}}\left[\mathrm{Var}(y\mid\boldsymbol{\beta},\boldsymbol{h})\right]+\mathrm{Var}_{\boldsymbol{\beta},\boldsymbol{h}\mid\boldsymbol{\lambda}_{\beta},\boldsymbol{\lambda}_{h}}\left[E(y\mid\boldsymbol{\beta},\boldsymbol{h})\right]$
$=E_{\boldsymbol{h}\mid\boldsymbol{\lambda}_{h}}[\text{exp}\{\boldsymbol{h}\}]+\mathrm{Var}_{\boldsymbol{\beta}|\lambda_{\beta}}[\boldsymbol{\beta}]$
$=\text{exp}\bigg{\{}\frac{1}{2}diag(\mathbf{S}\Lambda_{h}\mathbf{S}^{T})\bigg{\}}+\mathbf{S}\Lambda_{\beta}^{2}\mathbf{S}^{T}.$
The first term is diagonal, while the second term is dense due to the cumulative sum operation of $\mathbf{S}$.
The assumption of equal marginal distributions implies equality of variances:
$\text{exp}\bigg{\{}\frac{1}{2}diag(\mathbf{S}\Lambda_{h}\mathbf{S}^{T})\bigg{\}}+\mathbf{S}\Lambda_{\beta}\mathbf{S}^{T}=\text{exp}\bigg{\{}\frac{1}{2}diag(\mathbf{S}\Lambda_{h}^{*}\mathbf{S}^{T})\bigg{\}}+\mathbf{S}\Lambda_{\beta}^{*}\mathbf{S}^{T},$
$\text{exp}\bigg{\{}\frac{1}{2}diag(\mathbf{S}\Lambda_{h}\mathbf{S}^{T})\bigg{\}}-\text{exp}\bigg{\{}\frac{1}{2}diag(\mathbf{S}\Lambda_{h}^{*}\mathbf{S}^{T})\bigg{\}}=\mathbf{S}\Lambda_{\beta}^{*}\mathbf{S}^{T}-\mathbf{S}\Lambda_{\beta}\mathbf{S}^{T}.$
The left hand side is always diagonal while the right hand side is dense unless $\boldsymbol{\lambda}_{\beta}=\boldsymbol{\lambda}_{\beta}^{*}$. Hence, the only solution consistent with both sides is when both are zero matrix, implying $\boldsymbol{\lambda}_{\beta}=\boldsymbol{\lambda}_{\beta}^{*}$ and $\boldsymbol{\lambda}_{h}=\boldsymbol{\lambda}_{h}^{*}$. This is a contradiction to the assumption of distinct parameters: $(\boldsymbol{\lambda}_{\beta},\boldsymbol{\lambda}_{h})\neq(\boldsymbol{\lambda}_{\beta}^{*},\boldsymbol{\lambda}_{h}^{*}).$ Therefore, the parameters $(\boldsymbol{\lambda}_{\beta},\boldsymbol{\lambda}_{h})$ are identifiable.
Appendix B Full Conditional for Gibbs Sampling
B.1 $j$
$\boldsymbol{j}=(j_{1},\ldots,j_{T})$ was introduced to expand the likelihood on $\boldsymbol{y^{*}}:=(\log(y_{1}^{2}),\ldots,\log(y_{T}^{2}))$. In this section, we show that $\forall k\in\{1,\ldots,10\}$:
$\mathrm{p}(j_{t}=k|h_{t},v_{t},s_{t},\xi_{t},\mu,\xi_{\mu},\phi,y_{t}^{*})$ $=\mathrm{p}(j_{t}=k|h_{t},y_{t}^{*})$
$=\frac{\mathcal{N}(y_{t}^{*}|h_{t}+m_{k},w_{k}^{2})p_{k}}{\sum_{i=1}^{10}p_{i}\mathcal{N}(y_{t}^{*}|h_{t}+m_{i},w_{i}^{2})},$ $\forall t\in\{1,\ldots,T\}$
$\boldsymbol{j}$ is only associated with $y^{*}$, which is only associated with $\boldsymbol{h}$ and $\boldsymbol{j}$:
[…]
By Omori et al. [2007],
$f(y_{t}^{*}|j_{t},h_{t})=\mathcal{N}(y_{t}^{*}|h_{t}+m_{j_{t}},w_{j_{t}}^{2})$
Naturally, $\forall k\in\{1,\ldots,10\}$
$\mathrm{p}(j_{t}=k|h_{t},y_{t}^{*})$ $=\frac{f(y^{*}_{t}|h_{t},j_{t}=k)\mathrm{p}(j_{t}=k)}{f(y^{*}_{t}|h_{t})}$
$=\frac{f(y^{*}_{t}|h_{t},j_{t}=k)\mathrm{p}(j_{t}=k)}{\sum_{i=1}^{10}\mathrm{p}(j_{t}=i)f(y^{*}_{t}|h_{t},j_{t}=i)}$
$=\frac{\mathcal{N}(y_{t}^{*}|h_{t}+\mu_{k},\sigma_{k}^{2})p_{k}}{\sum_{i=1}^{10}p_{i}\mathcal{N}(y_{t}^{*}|h_{t}+\mu_{i},\sigma_{i}^{2})},$ $\forall t\in\{1,\ldots,T\},$
which is what we wanted to show. Exact distribution on $j_{t}\stackrel{{\scriptstyle i.i.d}}{{\sim}}Categorical(\pi^{\text{Omori}})$ as well as corresponding mean and the variance parameter of each component is described in Omori et al. [2007].
B.2 $h$
The likelihood and the conditional priors on $\boldsymbol{h}$ are:
$f(y^{*}|\boldsymbol{j},\boldsymbol{h})$ $=\prod_{t=1}^{T}\mathcal{N}(y_{t}^{*}|h_{t}+m_{j_{t}},w_{j_{t}}^{2})=\mathcal{N}(y^{*}|h+m_{j},w_{j}^{2}I)$
$f(\boldsymbol{h}|\textbf{v})$ $=\mathcal{N}(h_{1}|0,e^{v_{1}})\prod_{t=1}^{T}\mathcal{N}(h_{t}|h_{t-1},e^{v_{t}})$
Note that $\boldsymbol{h}$, a conditionally Gaussian, is a linear combination of $\Delta\boldsymbol{h}$, with the cumulative sum operator, the precision of the conditional prior of $h$ is a tridiagonal matrix $Q_{v}$
$Q_{v}:=\begin{bmatrix}\bigg{(}\frac{1}{e^{v_{2}}}+\frac{1}{e^{v_{1}}}\bigg{)}&-\frac{1}{e^{v_{2}}}&0&\ldots&\ldots&0\\ -\frac{1}{e^{v_{2}}}&\bigg{(}\frac{1}{e^{v_{3}}}+\frac{1}{e^{v_{2}}}\bigg{)}&-\frac{1}{e^{v_{3}}}&\ddots&\ddots&\vdots\\ 0&-\frac{1}{e^{v_{3}}}&\bigg{(}\frac{1}{e^{v_{4}}}+\frac{1}{e^{v_{3}}}\bigg{)}&-\frac{1}{e^{v_{4}}}&\ddots&\ddots\\ \vdots&\ddots&\ddots&\ddots&\ddots&0\\ \vdots&\ddots&\ddots&-\frac{1}{e^{v_{T-1}}}&\bigg{(}\frac{1}{e^{v_{T-1}}}+\frac{1}{e^{v_{T}}}\bigg{)}&-\frac{1}{e^{v_{T}}}\\ 0&\ldots&\ldots&0&-\frac{1}{e^{v_{T}}}&\frac{1}{e^{v_{T}}}\\ \end{bmatrix}$
Since both the likelihood and the prior are Gaussian, the conditional posterior is also Gaussian:
$f(\boldsymbol{h}|\boldsymbol{j},\boldsymbol{v},\boldsymbol{y^{*}})=\mathcal{N}\bigg{(}\boldsymbol{h}\bigg{|}\bigg{(}Q_{v}+I\frac{1}{w^{2}_{\boldsymbol{j}}}\bigg{)}^{-1}\frac{y-m_{\boldsymbol{j}}}{w^{2}_{\boldsymbol{j}}},\bigg{(}Q_{v}+I\frac{1}{w^{2}_{\boldsymbol{j}}}\bigg{)}^{-1}\bigg{)}.$
B.3 $v$
We use the all-without-loop (AWOL) sampler by Kastner and Frühwirth-Schnatter [2014] with parameter expansion on the error term with scale mixture normal distribution for sampling $\textbf{v},\mu$ and $\phi.$ By the conditional independence, the likelihood with respect to $\boldsymbol{v}$, reduces to $f(\boldsymbol{\omega^{*}}|v)$ where $\omega_{1}^{*}=\text{log}(h_{1}^{2})$ and $\omega_{t}^{*}=\text{log}((h_{t}-h_{t-1})^{2}),\forall t\geq 2.$ With the parameter expansion on the likelihood with 10-component Gaussian Mixture proposed by Omori et al. [2007],
$f(\boldsymbol{\omega}^{*}|\textbf{v},\boldsymbol{s})$ $=\mathcal{N}(\omega^{*}|v+m_{s},w_{s}^{2}I).$
The conditional prior distribution of $v$ is also Gaussian due to the Z-distribution being a scale mixture Gaussian. $\boldsymbol{v}$ is also a linear combination of $\boldsymbol{v}^{*}=(v^{*}_{1},\ldots,v^{*}_{T})^{\prime}$, where $v_{1}^{*}=v_{1}$ and $v^{*}_{t+1}=v_{t+1}-\phi v_{t},t\geq 2$, conditionally independent Gaussian random variables:
$f(v^{*}_{1}|\xi_{1},\mu)$ $=f(v_{1}|\xi_{0},\mu)=\mathcal{N}(v^{*}_{t}|\mu,\frac{1}{\xi_{1}})$
$f(v^{*}_{t}|\xi_{t-1},\mu,\phi)$ $=f(v_{t}-\phi v_{t-1}|\xi_{t-1},\mu,\phi)=\mathcal{N}(v^{*}_{t}|\mu(1-\phi),\frac{1}{\xi_{t-1}})$ $t\geq 2,$
Similar to section B.2, we have:
$f(\boldsymbol{v}|\boldsymbol{\xi},\mu,\phi)=\mathcal{N}(\boldsymbol{v}|0,Q_{\xi,\phi}^{-1}),$
with
$Q_{\xi,\phi}=\begin{bmatrix}\xi_{1}+\phi^{2}\xi_{2}&-\phi\xi_{2}&0&\ldots&\ldots&0\\ -\phi\xi_{2}&\xi_{2}+\phi^{2}\xi_{3}&-\phi\xi_{3}&\ddots&\ddots&\vdots\\ 0&-\phi\xi_{3}&\xi_{3}+\phi^{2}\xi_{4}&-\phi\xi_{4}&\ddots\\ \vdots&\ddots&\ddots&\ddots&\ddots&0\\ \vdots&\ddots&\ddots&-\phi\xi_{T-1}&\xi_{T-1}+\phi^{2}\xi_{T}&-\phi\xi_{T}\\ 0&\ldots&\ldots&0&-\phi\xi_{T}&\xi_{T}\\ \end{bmatrix}$
Thus,
$f(\textbf{v}|y^{*},\ldots)$ $=\mathcal{N}\bigg{(}\textbf{v}\bigg{|}\bigg{(}Q_{\xi,\phi}+I\frac{1}{w^{2}_{\boldsymbol{s}}}\bigg{)}^{-1}\bigg{(}\frac{\boldsymbol{\omega^{*}}-m_{\boldsymbol{s}}}{w^{2}_{\boldsymbol{s}}}+Q_{\xi,\phi}\boldsymbol{1}\mu\bigg{)},\bigg{(}Q_{\xi,\phi}+I\frac{1}{w^{2}_{\boldsymbol{s}}}\bigg{)}^{-1}\bigg{)}$
B.4 $s$
In section B.2, $\boldsymbol{s}=(s_{1},\ldots,s_{T})^{\prime}$ was introduced to expand $\boldsymbol{\omega}^{*}|\textbf{v}$. Based on the same argument used in section B.1, $\forall k\in\{1,\ldots,10\}$
$\mathrm{p}(s_{t}=k|\omega_{t}^{*},v_{t})$ $=\frac{\mathcal{N}(\omega_{t}^{*}|v_{t}+m_{k},w_{k}^{2})p_{k}}{\sum_{i=1}^{10}p_{i}\mathcal{N}(\omega^{*}_{t}|v_{t}+m_{i},w_{i}^{2})},$ $\forall t\in\{1,\ldots,T\}.$
B.5 $\xi$
Based on Polson et al. [2013], we have
$v_{1}$ $=\mu+\eta_{0}$
$v_{t}$ $=\mu+\phi(v_{t-1}-\mu)+\eta_{t-1},$ $\forall t\geq 2.$
Naturally, $\eta_{1}=v_{1}-\mu$ and $\eta_{t}=v_{t+1}-\phi v_{t}-\mu(1-\phi),\forall t\geq 2$.
$f(\xi_{0}|\textbf{v},\mu,\phi)$ $=\mathcal{PG}(\xi_{0}|1,v_{1}-\mu)$
$f(\xi_{t-1}|\textbf{v},\mu,\phi)$ $=\mathcal{PG}(\xi_{t-1}|1,v_{t}-\phi v_{t-1}-\mu(1-\phi))$ $\forall t\geq 2,$
where $\mathcal{PG}$ represents the density function for Polya-Gamma random variable.
B.6 $\mu$
Define $\hat{v}_{\phi}^{*}=\frac{\sum_{t=2}^{T}\sqrt{\xi_{t}}(v_{t}-\phi v_{t-1})}{(1-\phi)\sum_{t=2}^{T}\sqrt{\xi_{t}}}$, so that
$v_{t}-\phi v_{t-1}|\boldsymbol{\xi},\mu,\phi\sim N((1-\phi)\mu,\frac{1}{\xi_{t}})$
$\hat{v}_{\phi}^{*}|\boldsymbol{\xi},\mu,\phi\sim N(\mu,\sigma^{2}_{\xi,\phi})$
$\sigma^{2}_{\xi,\phi}=\frac{T-1}{((1-\phi)\sum_{t=2}^{T}\sqrt{\xi_{t}})^{2}}.$
The conditional prior distribution on $\mu$ also reduces to $f(\mu|\xi_{\mu})=\mathcal{N}(\mu|0,\frac{1}{\xi_{\mu}})$ Thus, the conditional posterior distribution for $\mu$ is
$f(\mu|\hat{v}^{*}_{\phi},\boldsymbol{\xi},\xi_{\mu},\phi)$ $=\mathcal{N}\bigg{(}\mu\bigg{|}\bigg{(}\frac{1}{\sigma^{2}_{\xi,\phi}}+\xi_{\mu}\bigg{)}^{-1}\frac{\hat{v}^{*}_{\phi}}{\sigma^{2}_{\xi,\phi}},\bigg{(}\frac{1}{\sigma^{2}_{\xi,\phi}}+\xi_{\mu}\bigg{)}^{-1}\bigg{)}.$
B.7 $\xi_{\mu}$
Based on the scale mixture representation of the Z-distribution Polson et al. [2013]:
$f(\xi_{\mu}|\mu)=\mathcal{PG}(\xi_{\mu}|1,\mu).$
B.8 $\phi$
Unlike other parameters, $\phi$ is not conditionally conjugate to its likelihood. Thus, the slice sampling method Neal [2003] is used for the MCMC sampling. By taking the linear transformation of $v_{t}$, Let’s consider its likelihood becomes $\hat{v}_{\mu}:=\frac{1}{T-1}\sum_{t=2}^{T}v_{\mu,t}$, where $v_{\mu,t}=\frac{1}{2}\bigg{(}\frac{v_{t}-\mu}{v_{t-1}-\mu}+1\bigg{)}$. We have the following conditional likelihood and prior for $\phi$:
$f(\hat{v}_{\mu}|\mu,\boldsymbol{\xi})=\mathcal{N}\bigg{(}\hat{v}_{\mu}\bigg{|}\frac{\phi+1}{2},\frac{1}{(T-1)^{2}}\sum_{t=2}^{T}\frac{1}{4\xi_{t-1}(v_{t-1}-\mu)^{2}}\bigg{)}$
$f(\frac{\phi+1}{2})\sim Beta\bigg{(}\frac{\phi+1}{2}\bigg{|}10,2\bigg{)}$